OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. x + 5 = x + 5 A. Reflexive Property B. Symmetric Property C. Associative Property of Addition D. Distributive Property

jimthompson5910 (jim_thompson5910):

The symmetric property says that if x = y, then y = x Basically you can flip the equation and it won't change things Is that what's going on here?

OpenStudy (anonymous):

YEAH i think so.

jimthompson5910 (jim_thompson5910):

Well it may look like it happens, there's actually a more better answer (that's more accurate of what is going on)

OpenStudy (anonymous):

Is it kind of like this same question?? Name the algebraic property demonstrated in the example below. 3 • (x • y) = (3 • x) • y A. Distributive Property B. Transitive Property C. Associative Property of Multiplication D. Commutative Property of Multiplication

jimthompson5910 (jim_thompson5910):

If I say 2 = 2 (sounds silly, but it's important), then this is definitely true and it's true because any number is equal to itself due to the reflexive property.

OpenStudy (anonymous):

The original question is actually the commutative property of addition

jimthompson5910 (jim_thompson5910):

It's only commutative if you say x + 5 = 5 + x

OpenStudy (anonymous):

oh i read it wrong

OpenStudy (anonymous):

so for the first question what would be the correct answer?

jimthompson5910 (jim_thompson5910):

x is a number 5 is a number so x+5 is a number When we say x+5 = x+5 we're basically saying SOME NUMBER = ITSELF and we can say this because ???

OpenStudy (anonymous):

communitive

jimthompson5910 (jim_thompson5910):

no, that property allows to say x+5 = 5+x ie, you can add in any order

jimthompson5910 (jim_thompson5910):

allows us*

OpenStudy (anonymous):

so symmetric property

jimthompson5910 (jim_thompson5910):

that is the property that says if x = y then y = x It's close to what we have, but there's actually a more accurate answer